# A Single-Step Non-Autoregressive Automatic Speech Recognition Architecture with High Accuracy and Inference Speed

### v1

@@ 3.2 Results on AISHELL-1 and AISHELL-2
The evaluation results of AISHELL-1 and AISHELL-2 are detailed in Table 1. CER and RTF of previous work are obtained from the paper’s report, except that the results of ESPnet Conformer are obtained from the ESPnet official repository. The RTF of our models is calculated on the AISHELL-1 test set with a batch size of $1$. For a fair comparison with the published work, none of our experiments in Table 1 uses an external LM or pre-training.
@@
For AISHELL-1 task, the proposed EfficientASR models (both base and large size) outperform all other models presented. On the AISEHLL-1 test set, our large model achieves a CER of 4.62%, which is 0.58% (absolutely) better than the previous best result of NAR models (from Paraformer [24]). Furthermore, the inference speed of our large model is twice more than that of Paraformer.
@@
When compared with the AR models, our model not only outperforms SOTA AR Conformer in CER, but also has inference speed 30x faster than AR Conformer, which is an impressive performance. Since AISHELL-1 contains less than 200 hours traning data, we also validate our method on a larger corpus, 1000 hours AISHELL-2. On the AISHELL-2 test ios set, our large model achieves a CER of 5.76%, which also outperforms all AR and NAR models presented.
3.3 Analysis
To more fairly evaluate the performance of our model, we use ESPnet[29] to reproduce the SOTA AR and NAR models, namely Conformer and Paraformer. The results are shown in Table 2. With a similar model size, the base version of EfficientASR achieves the lowest CER and RTF in the Aishell-1 test set, which verifies the effectiveness of the proposed method.
Table 2: Reproduction CER (%) and RTF results of the SOTA AR model Conformer and NAR model Paraformer on the AISHELL-1 task and comparison with our model.
Model AR/NAR dev test RTF ($\downarrow$)
Conformer AR 4.30 4.70 0.3980
Paraformer NAR 4.73 5.11 0.0136
EfficientASR Base NAR 4.30 4.66 0.0058
Table 3: Reproduction Inference time(s) in different modules of Paraformer and EfficientASR Base. Each item is the cumulative inference time on the entire AISHELL-1 test set.
Model Encoder Predictor Decoder
Paraformer 208.76 191.79 56.58
EfficientASR Base 209.55 9.65 32.60
In Table 3, we give a detailed analysis of inference time in different modules of Paraformer and the base version of EfficientASR. Note that here Paraformer’s predictor includes the CIF module, while ours includes the attention reconstruction module. As can been seen, the EfficientASR Base is much faster in the predictor module. This is mainly caused by the CIF mechanism in Paraformer. When decoding, CIF needs to recursively obtain the acoustic boundary of each emitted token in advance, while our model performs parallel calculations during the entire inference process. In addition, our decoder is also faster. This is because Paraformer’s decoder is built using transformer decoder blocks, while our decoder is built using transformer encoder blocks that do not include cross-attention layers.
Table 4: Analysis of decoding results (%). Baseline: the alignment for attention reconstruction is obtained from the predictor. Oracle: the alignment for attention reconstruction is obtained from generator based on ground-truth labels. Note: Using only the Baseline decoding method, we have achieved SOTA CER performance on corresponding benchmarks.
AISHELL-1 AISHELL-2
dev test dev-ios test-ios
Baseline (SER) 2.12 2.43 3.09 3.29
Baseline (CER) 4.26 4.62 5.64 5.76
   + 4-gram 4.06 4.37 5.31 5.40
Oracle (CER) 3.61 3.93 3.78 4.10
In Table 4, we further explore the potential of the proposed model. Chinese characters are monosyllabic characters in which different characters may have the same Pinyin syllable (the phonetic symbols of Mandarin). When we analyze the decoding results, we find that many misrecognized characters have the same syllable as their reference characters, in other words, although these characters are misrecognized, their syllables are correctly recognized. So we use a pronunciation lexicon to map the character-based decoding results to syllable-based decoding results to test the syllable error rate (SER) (e.g. Chinese characters ”{CJK*}UTF8gbsn你 好 ”, and their syllables are ”$ni3\ hao3$”). As is shown in Table 4, SER is significantly lower than CER. This means that if we can eliminate this kind of errors, there is still a lot of room for improvement in CER performance. Therefore, we consider using beam search decoding (the beam size is $10$) combined with a 4-gram LM to reduce such errors. The 4-gram LMs are trained using the training set text of the corresponding corpus. It can be seen that the CER results using 4-gram have been improved to a certain extent. However, there is still a large gap between CER and SER, which means that the proposed EfficientASR still has great potential for improvement in semantic modeling. This provides guidance for our future work.
On the other hand, to evaluate the modeling ability of the alignment predictor, we analyze the gap between Baseline CER and Oracle CER, where Baseline’s alignment is obtained from the alignment predictor and the Oracle’s alignment is obtained from the alignment generator based on the ground-truth labels. As shown in Table 4, the Oracle CER is much lower than the Baseline CER. Since the alignment predictor in our model is decoupled from the generator, the performance of the model can be improved to a new level if the predictor can better learn the alignment produced by the generator, which provides another direction for the optimization of the model.
Figure 3: Alignment plots of both the baseline model and proposed model. The first plot is the alignment plot of the Conformer, while the second plot is the alignment constructed by the proposed EfficientASR. The horizontal axis represents the input frame step, and the vertical axis represents the output step.
@@ 3.3 Analysis
To further analyze the behaviour of the proposed model, we plot the alignment matrices comparison between the Conformer and the proposed EfficientASR in Figure 3. The alignment matrix of Conformer is from the last cross-attention layer of the model. As can be seen, the proposed method generates more monotonic and cleaner alignments compared to Conformer.
4 Conclusion
@@
In this paper, we propose a novel single-step NAR ASR architecture called EfficientASR. By leveraging our proposed methods for generating monotonic alignments and constructing alignment matrices, EfficientASR not only possesses a simpler structure but also achieves leading performance in terms of inference efficiency and accuracy. Experiments on the open source dataset AISHELL-1 demonstrate that the proposed EfficientASR achieves comparable CER performance to SOTA AR Conformer, with a 30x decoding speedup. On the other hand, we find that the CER performance of our model has great potential for improvement. Our next work will focus on improving the semantic modeling ability of the proposed EfficientASR. Besides, we will study the performance of the EfficientASR in alphabetic languages (e.g. English).

### v4

@@ III-B Results
The evaluation results of AISHELL-1 and AISHELL-2 are detailed in Table I. The CER and RTF values of previous works are obtained from the papers’ report, except that results of the ESPnet Conformer are obtained from the ESPnet official repository. The RTF of our models is calculated on the AISHELL-1 test set with a batch size of $1$. For a fair comparison with the published work, none of our experiments in Table I uses an external language model (LM) or pre-training.
@@
For AISHELL-1 task, the proposed EffectiveASR models (both base and large size) outperform all other models presented. On the AISEHLL-1 test set, our large model achieves a CER of 4.62%, which is 0.58%/0.17% (absolutely) better than the previous leading results of NAR models (from Paraformer [29]/E-Paraformer [24]). Furthermore, the inference speed of our large model is twice more than that of Paraformer. With the application of PIF, E-Paraformer also outperforms Paraformer in both inference speed and CER. However, the two-pass training process of E-Paraformer makes the model less compact and less efficient compared to the single-step architecture of EffectiveASR. When compared with the AR models, our model not only outperforms AR Conformer in CER, but also has inference speed 30x faster than the AR Conformer, which is an impressive performance.
@@
Since AISHELL-1 contains less than 200 hours traning data, we also validate our method on a larger corpus, 1000 hours AISHELL-2. On the AISHELL-2 test ios set, our large model achieves a CER of 5.76%, which also outperforms all AR and NAR models presented.
@@
To further analyze the behavior of the proposed model, we plot the alignment matrices in Figure 3, comparing the Conformer with the proposed EffectiveASR. The alignment matrix of the Conformer is taken from the model’s final cross-attention layer. As depicted in the figure, the proposed approach yields more consistent and refined alignments compared to the Conformer model, thereby contributing to an enhanced CER performance.
Figure 3: Alignment plots of both the baseline and proposed model. The first plot is the alignment plot of the Conformer, while the second plot is the alignment constructed by the proposed EffectiveASR. The horizontal axis represents the input frame step, and the vertical axis represents the output step.
IV Conclusion
@@
In this paper, we propose a novel single-step NAR ASR architecture called EffectiveASR. By leveraging our proposed methods for generating monotonic alignments and constructing alignment matrices, EffectiveASR not only possesses a simpler structure but also achieves leading performance in terms of inference efficiency and accuracy. Experiments on open-source Mandarin datasets demonstrate that the proposed EffectiveASR achieves better CER performance to the leading AR Conformer, with a 30x decoding speedup. In the future, we will focus on exploring the application of EffectiveASR in English speech recognition.